OpenStudy (anonymous):

Help please Is this function below is symmetric to the x-axis, y-axis, orign, or has has no symmetry at all.

OpenStudy (anonymous):

y = - 4x ^3

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large f(-x) = -4(-x)^3 = 4x^3 \ne f(x)\) so the function is NOT symmetric about `y` axis

ganeshie8 (ganeshie8):

still we need to check whether it is symmetric about x axis or origin..

ganeshie8 (ganeshie8):

know how to check them ?

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

To check symmetry about x axis : replace `y` with `-y` if the function does not change, then it is symmetrical about x axis

ganeshie8 (ganeshie8):

\(\large y = -4x^3\) replace `y` by `-y`, you get : \(\large -y = -4x^3 \) \(\large \implies y = 4x^3 \ne f(x)\) So the function is NOT symmetric about x axis also

OpenStudy (anonymous):

ok what would be 

ganeshie8 (ganeshie8):

are we done with the first question ? :)

ganeshie8 (ganeshie8):

we still need to check if it has symmetry about `origin` right ?

OpenStudy (anonymous):

oh yea sorry

ganeshie8 (ganeshie8):

its okay... try below for testing symmetry about `origin` : y = - 4x ^3 replace `x with -x` and `y with -y` if the function does not change, then it is symmetrical about `origin`

ganeshie8 (ganeshie8):

y = - 4x ^3 replace `x with -x` and `y with -y` what do u get ?

OpenStudy (anonymous):

what do i look for i dont understand

ganeshie8 (ganeshie8):

y = -4x^3 replace x with -x and y with -y you get : -y = -4(-x)^3

ganeshie8 (ganeshie8):

if you simplify, you will see that its exactly same as the original function : -y = -4(-x)^3 -y = 4x^3 y = -4x^3 So the function is symmetric about the origin

ganeshie8 (ganeshie8):

if that makes more or less sense...